Suppose an economy has $200,000 of demand deposits and $40,000 of excess reserves, with a 10% required reserve ratio. If the monetary authorities lower the required reserve ratio to 2%: A. the excess reserves will fall by 10%. B. the excess reserves will rise. C. excess reserves will decrease by $20,000. D. there will be no more excess reserves in the system.
Answer #1

excess reserves at 10%=40000

required reserves at 10%=200000*10%=20000

actual reserves=excess reserves at 10%+required reserves at 10%

=40000+20000=60000

required reserves at 2%=200000*2%=4000

excess reserves at 2%=actual reserves-required reserves at 2%

=60000-4000

=56000

so excess reserves has increased from 40000 to 56000, so answer is

the excess reserves will rise